# Superuniversal Statistics with Topological Origins for non-Hermitian Scattering Singularities

### v1

@@ I.3 Paper Overview
This paper is structured as follows. Section II describes the experimental systems considered and their parametric degrees of freedom, as well as the measurements. In section III, the experimental data is presented and discussed in three subsections: III.1 has the statistics of fundamental singularities, III.2 presents the absorption dependence of singularity abundance, and III.3 contains the statistics of higher order singularities, which are composite of multiple fundamental singularities. There are also appendices A-F which document details and additional data supporting the claims made in the text.
Although the experimental data presented in this paper all comes from microwave systems, the agreement shown with Random Matrix Theory (RMT) formulation of the scattering matrix indicates that the conclusions are generic for all wave scattering systems. We suggest that this is not the limit, and the “superuniversal” statistics found in these systems extend to many other fields of physics where similar analysis can be performed.
@@ II Experiment
We measure the $M\times M$ $S$-matrix of various complex microwave scattering systems with calibrated Keysight PNA-X N5242A and PNA-X N5242B microwave vector network analyzers. By complex we mean that the systems are excited with waves whose wavelengths are small compared to the system size, and interference effects for waves following different ray trajectories are extremely sensitive to details of the system configuration. The systems in question are depicted in Fig. 2: (a) a tetrahedral graph ($\mathcal{D}=1$), (b) a chaotic quarter bowtie billiard ($\mathcal{D}=2$), and (c) a three dimensional cavity with various symmetry breaking elements ($\mathcal{D}=3$), where $\mathcal{D}$ is the wave propagation dimension. For clarity in the schematic, the $M=2$ scattering channels connected to the PNA are shown in red, and the embedded perturbers are shown in green. In the graph, the perturbers take the form of voltage controlled mechanical phase shifters on four of the six bonds [107, 108, 109, 110, 111]. These phase shifters act as variable-length cables $L+\delta L$, allowing us to change the interference conditions at the nodes. Mechanical phase shifters are superior to digital phase shifters for this purpose, as digital phase shifters typically have much higher insertion loss, whereas the mechanical phase shifters have loss comparable to the coaxial cables that make up the rest of the graph, and have very precise step size resolution. In the billiard and three dimensional cavity there are globally-biased, voltage controlled varactor-loaded metasurfaces [112, 113, 114, 115, 59]. The metasurfaces give us control over the amplitude and phase of reflected waves, and in the reverberant scattering environments most waves will interact with these metasurfaces multiple times, allowing for impactful control over the scattering properties of the systems.
@@
Figure 2: Schematics of experimental systems. (a) A tetrahedral microwave graph ($\mathcal{D}=1$), (b) a ray-chaotic quarter bowtie billiard ($\mathcal{D}=2$), and (c) a three dimensional cavity with various symmetry breaking elements ($\mathcal{D}=3$). Scattering channels connected to the PNA network analyzer (top right) are marked in red, and embedded perturbers are marked in green.
@@
The tunable perturbers make two types of measurements possible. One is a continuous parameter sweep like the kind shown in Fig. 1 where typically a single perturber is tuned with very fine steps between frequency scans, while the rest of the system parameters are held fixed. This kind of measurement reveals the phase winding around singularities, and through the use of a second perturber the dynamics of the singularities can be visualized [59, 91]. The other kind of measurement is a statistical ensemble, where each perturber is tuned simultaneously to very different values between frequency scans. This ensures that the difference between each realization in the ensemble is statistically significant. In other words, the first kind of measurement gives the scattering matrix in a closed area ($d=2$) or volume ($d=3$) of parameter space, while the second kind of measurement provides a reasonably ergodic sampling of a higher dimensional parameter space ($d\geq 4$). All the statistical results discussed in this paper will be derived from the second kind of measurement.
@@
We consider four global system parameters as being important and hold these fixed across realizations of a given ensemble. These are: the dimension for wave propagation $\mathcal{D}$ ($\mathcal{D}\in\{1,2,3\}$), number of scattering channels $M$ ($M\in\{1,\dots,\infty\}$, although our PNA has a limit of $M\leq 4$), Dyson class $\beta$ ($\beta\in\{1,2\}$), and uniform dissipation $\eta$ ($\eta\in\{0,\dots,\infty\}$). The Lorentz reciprocity of a system can be broken, changing $\beta$ from 1 to 2, by adding a magnetized ferrite in the propagation path of the microwaves [116, 117, 118, 119, 120, 121, 122]. Note that $\eta$ is a dimensionless quantity given by $\eta:=\tau_{H}\tilde{\eta}$ where $\tilde{\eta}$ is an absorption rate proportional to the average mode bandwidth [123, 124, 34]. The Heisenberg time $\tau_{H}=\frac{2\pi}{\Delta}$ can be related to the mean dwell time of waves in a scattering environment, where $\Delta$ is the mean mode spacing in frequency units. Hence, the absorption strength $\eta$ is inherently tied to our system dimension $\mathcal{D}$, as the larger systems have longer dwell times which cause the waves to be more strongly attenuated. Since both $\tilde{\eta}$ and $\tau_{H}$ (for $\mathcal{D}>1$) are frequency dependent [43], by measuring in different frequency bands we can systematically vary the value of $\eta$. Distributing absorbers in a cavity [125, 126, 127, 128] or attenuators on the bonds of a graph [36, 111, 129] is another way to change $\eta$, though these methods can only increase the loss. One can also employ cryogenics to decrease $\eta$ [127, 130, 44]. For this paper we have prepared systems with $\eta$ varying from $1.8$ to $145$, which covers well what has been considered the low, moderate, and high absorption strength regimes [131, 132, 125, 133, 134, 135, 129].
@@
The $S$-matrix measured by a PNA is dependent on how well the scattering channels are coupled to the enclosure. Our systems are designed to be well coupled, but the coupling is still imperfect and frequency dependent. We remove these non-universal effects from ensembles through application of the Random Coupling Model (RCM) normalization process, which results in $S$-matrices with perfect coupling [125, 136, 137, 138, 139, 140, 141, 142, 143, 144, 145, 108, 146]. Coupling effects cannot be removed from parametric studies like in Fig. 1 because RCM requires statistical information. The effects of coupling on scattering singularity statistics are not addressed in this paper and will be explored in subsequent work. For more details on the experimental systems, such as more information on the tunable pertubers or how an ensemble value of $\eta$ is calculated, refer to Refs. [53] and [91].
@@ III Data and Discussion
In this section we present statistical results of scattering singularities from both experiment and RMT numerics. Note that every scattering matrix included is perfectly coupled, see Section II. We consider PDFs of functions that diverge at singularities for two reasons: (1) divergent functions are more advantageous for finding singularities, as compared to complex linear functions, because they require only one degree of freedom and have a more telling character when singularities occur, and (2) the PDF of a divergent function is dominated by an infinitely long tail which corresponds to the presence of singularities. This means most of the PDF is of importance when considering singularities, rather than just a small region around the singularity value. In particular, for every fundamental singularity, it is possible to define a time delay or some generalized delay that diverges at the singularity, as discussed in Appendix. E. In subsection III.1 we present the statistics of topologically protected points (fundamental singularities) and curves in two dimensional parameter space. Subsection III.2 focuses on how absorption affects singularity abundance in generic systems because absorption appears to be the most important determining parameter. Lastly, in subsection III.3 we discuss higher order singularities which show system-specific statistics, in contrast to fundamental singularities.
@@ III.1 Fundamental Singularities - Statistics
Figure 3: (a-b) PDFs of $\frac{\langle|\text{Re}[\delta R_{xy}]|\rangle}{|\text{Re}[\delta R_{xy}]|}$ and $\frac{\langle|\delta R_{xy}|\rangle}{|\delta R_{xy}|}$ from 31 ensembles of experimental scattering systems with different values for the four parameters $\mathcal{D},M,\beta,\eta$. Color gradient of curves corresponds to $\eta$ (not uniformly spaced) of the ensembles. Dashed black lines characterize power law of tail behavior. (c-d) PDFs of $\frac{\langle|\text{Re}[\mathcal{S}]|\rangle}{|\text{Re}[\mathcal{S}]|}$ and $\frac{\langle|\mathcal{S}|\rangle}{|\mathcal{S}|}$ for various functions $\mathcal{S}$ from two arbitrary ensembles. Dashed black lines characterize power law of tail behavior.
@@
In Fig. 3(a-b), we show the logarithmic PDFs of $\frac{\langle|\text{Re}[\delta R_{xy}]|\rangle}{|\text{Re}[\delta R_{xy}]|}$ and $\frac{\langle|\delta R_{xy}|\rangle}{|\delta R_{xy}|}$ from 31 ensembles of experimental scattering systems with different values for the four parameters $\mathcal{D},M,\beta,\eta$, where $\delta R_{xy}=S_{xx}-S_{yy}$ is the reflection difference between ports $x$ and $y$, and $\langle|\delta R_{xy}|\rangle$ is the ensemble mean of $|\delta R_{xy}|$. We normalize by the ensemble mean value to bring the PDFs together, as otherwise the PDFs from different systems can be distantly shifted horizontally from each other, making it difficult to compare the tails. We find that $\mathcal{P}(\frac{\langle|\text{Re}[\delta R_{xy}]|\rangle}{|\text{Re}[\delta R_{xy}|]})$ has a universal $-2$ power law tail that holds for all systems, represented by the dashed black line in panel (a), since $\frac{\langle|\text{Re}[\delta R_{xy}]|\rangle}{|\text{Re}[\delta R_{xy}]|}$ diverges at $\text{Re}[\delta R_{xy}]=0$, a constraint on a single degree of freedom. Similarly, $\frac{\langle|\delta R_{xy}|\rangle}{|\delta R_{xy}|}$ diverges at $\delta R_{xy}=0+i0$ which are points, specifically symmetric reflection points, and we find that $\mathcal{P}(\frac{\langle|\delta R_{xy}|\rangle}{|\delta R_{xy}|})$ has a universal $-3$ power law represented by the dashed line in panel (b).
@@
Three arbitrary choices were made in constructing Fig. 3(a-b). First, the choice of complex scalar function to investigate. There is nothing special about the reflection difference $\delta R_{xy}$, any function with singularities will produce the same statistical results. Second, in panel (a) we could have used the imaginary part of $\delta R_{xy}$ (as opposed to the real part) and the PDF tail behavior would have been the same; there is no particular reason to chose one over the other. Finally, plotting the PDFs of the magnitude, without regard to sign. This choice was made for simplicity as plotting both positive and negative values on a logarithmic scale is non-trivial. That was done, however, in Refs. [52, 53] for the case of complex time delay, and there the same tails were seen on both the positive and negative sides of the distribution.
@@
In Fig. 3(c), we show the logarithmic PDFs of eight different functions which diverge at topologically protected curves: (i) $\frac{\langle|C|\rangle}{|C|}$, (ii) $\frac{\langle|\text{Re[det}S]|\rangle}{|\text{Re[det}S]|}$, (iii) $\frac{\langle|\text{Re}[S_{xx}]|\rangle}{|\text{Re}[S_{xx}]|}$, (iv) $\frac{\langle|\text{Re}[S_{xy}]|\rangle}{|\text{Re}[S_{xy}]|}$, (v) $\frac{\langle|\text{Re}[\delta R_{xy}]|\rangle}{|\text{Re}[\delta R_{xy}]|}$, (vi) $\frac{\langle|\text{Re}[\tau_{xx}]|\rangle}{|\text{Re}[\tau_{xx}]|}$, (vii) $\frac{\langle|\text{Re}[\tau_{xy}]|\rangle}{|\text{Re}[\tau_{xy}]|}$, (viii) $\frac{\langle|\text{Re}[\tau_{WS}]|\rangle}{|\text{Re}[\tau_{WS}]|}$. Just as in panel (a), the same thing could be done using the PDFs of the imaginary parts of these functions rather than the real, except for $\frac{\langle|C|\rangle}{|C|}$ as $|C|$ is already a real number. All the data in panel (c) comes from one ensemble, a reciprocal two channel billiard ($\beta=1,M=2,\mathcal{D}=2$) with uniform absorption of $\eta=40$. Because the functions diverge at curves, the PDFs all have $-2$ power laws as represented by the dashed black curve. The deepest blue curve corresponding to $\mathcal{P}(\frac{\langle|C|\rangle}{|C|})$ has a down turn and transitions to a $-3$ power law. This can be explained by the fact that our experimental systems always have some small degree of non-reciprocity which is impossible to eliminate. Further, the PNA determines $S_{xy}$ and $S_{yx}$ independently and does not have arbitrary precision. For most purposes, this is a negligible effect but in non-reciprocal systems two-eigenvector orthogonality takes the dimension of points, which induces the $-3$ power law. In RMT simulations that are perfectly reciprocal, this transition in the power law of the PDF tail from $-2$ to $-3$ is not seen.
@@
Panel (d) of Fig. 3 has the logarithmic PDFs of ten different functions which diverge at topologically protected points: (i) $\frac{\langle|C|\rangle}{|C|}$ diverges at Orthogonality Points, (ii) $\frac{1-\langle|C|\rangle}{1-|C|}$ diverges at Exceptional Points, (iii) $\frac{\langle|\text{det}S|\rangle}{|\text{det}S|}$ diverges at Coherent Perfect Absortion, (iv) $\frac{\langle|S_{xx}|\rangle}{|S_{xx}|}$ diverges at Reflectionless Scattering Modes, (v) $\frac{\langle|S_{xy}|\rangle}{|S_{xy}|}$ diverges at Transmissionless Scattering Modes, (vi) $\frac{\langle|\delta R_{xy}|\rangle}{|\delta R_{xy}|}$ diverges at Symmetric Reflection Points, (vii) $\frac{\langle|\delta T_{xy}|\rangle}{|\delta T_{xy}|}$ diverges at Reciprocal Transmission Points, (viii) $\frac{\langle|\tau_{xx}|\rangle}{|\tau_{xx}|}$ diverges at Complex Reflection Time Delay zeros, (ix) $\frac{\langle|\tau_{xy}|\rangle}{|\tau_{xy}|}$ diverges at Complex Transmission Time Delay zeros, (x) $\frac{\langle|\tau_{WS}|\rangle}{|\tau_{WS}|}$ diverges at Complex Wigner-Smith Time Delay zeros. All the data in this panel comes from an ensemble measurement of a non-reciprocal two channel graph ($\beta=2,M=2,\mathcal{D}=1$) with uniform absorption of $\eta=8$. Eigenvector orthogonality has dimension of $d-2$ in non-reciprocal systems, but dimension of $d-1$ in reciprocal systems, which is why $\mathcal{P}(\frac{\langle|C|\rangle}{|C|})$ appears in both panels (c) and (d). Unlike panel (c), there is no deviation from the expected $-3$ power law behavior for the tails in panel (d).
@@
Between panels (a) and (c), it should become clear that the PDF of every function in every system that diverges at a topologically protected curve has a $-2$ power law. Similarly, panels (b) and (d) convey that the PDF of every function in every system that diverges at a topologically protected point has a $-3$ power law. A very simple mathematical argument for why the universal power laws take these values is given in Appendix A. A figure showing the power laws of PDFs calculated from $S$-matrices constructed with RMT numerics is available in Appendix C. The fact that RMT numerics returns statistics for scattering singularities which is identical to our empirical results from microwave systems reveals that these statistics are not specific to any kind of wave scattering. Rather, these statistical results describe all generic scattering systems.
@@
Since a distribution with a $-3$ power law has an undefined variance, a generic non-Hermitian scattering system must have every kind of scattering singularity allowed somewhere in its parameter space. However, it may not necessarily be accessible depending on the tunability of the experimental perturbers. Note also that since the $-3$ power law tail is superuniversal, the PDF of a function that diverges at a singularity is the perfect platform for investigating which global system parameters have the greatest effect on singularity abundance and density. It is also worth considering what the limitations of the superuniversality are, and whether there are ways to break it, for example by having imperfect coupling which was seen to disrupt the $-3$ power laws from some but not all complex time delay PDFs in Ref. [53].

### v3

@@ I.3 Paper Overview
This paper is structured as follows. Section II describes the experimental systems considered and their parametric degrees of freedom, as well as the measurements. In section III, the experimental data is presented and discussed in three subsections: III.1 has the statistics of fundamental singularities, III.2 presents the absorption dependence of singularity abundance, and III.3 contains the statistics of higher order singularities, which are composite of multiple fundamental singularities. There are also appendices A-H which document details and additional data supporting the claims made in the text.
Although the experimental data presented in this paper all comes from microwave systems, the agreement shown with Random Matrix Theory (RMT) formulation of the scattering matrix indicates that the conclusions are generic for all wave scattering systems. We suggest that this is not the limit, and the “superuniversal” statistics found in these systems extend to many other fields of physics where similar analysis can be performed , such as the statistics of transmission zeros in optics or singularities of quantum weak measurements as seen in Ref. [Solli2004].
@@ II Experiment
We measure the $M\times M$ $S$-matrix of various complex microwave scattering systems with calibrated Keysight PNA-X N5242A and PNA-X N5242B microwave vector network analyzers. By complex we mean that the systems are excited with waves whose wavelengths are small compared to the system size, and interference effects for waves following different ray trajectories are extremely sensitive to details of the system configuration. The systems in question are depicted in Fig. 2: (a) a tetrahedral graph ($\mathcal{D}=1$), (b) a ray-chaotic quarter bowtie billiard ($\mathcal{D}=2$), and (c) a three dimensional cavity with various symmetry breaking elements ($\mathcal{D}=3$), where $\mathcal{D}$ is the wave propagation dimension. For clarity in the schematic, the $M=2$ scattering channels connected to the network analyzer are shown in red, and the embedded tunable perturbers are shown in green. In the graph, the tunable perturbers take the form of voltage controlled mechanical phase shifters on four of the six bonds [Rehemanjiang2018, Dietz2017, Che2021, Che2022, Lawniczak2023]. These phase shifters act as variable-length cables $L+\delta L$, allowing us to change the interference conditions at the nodes. Mechanical phase shifters are superior to digital phase shifters for this purpose, as digital phase shifters typically have much higher insertion loss, whereas the mechanical phase shifters have loss comparable to the coaxial cables that make up the rest of the graph, and have very precise step size resolution. In the billiard and three dimensional cavity there are globally-biased, voltage controlled varactor-loaded metasurfaces [Chen2016, Elsawy2023, Frazier2022, Sleasman2023, Erb2024]. The metasurfaces give us control over the amplitude and phase of reflected waves, and in the reverberant scattering environments most waves will interact with these metasurfaces multiple times, allowing for impactful control over the scattering properties of the systems.
@@
Figure 2: Schematics of experimental systems. (a) A tetrahedral microwave graph ($\mathcal{D}=1$), (b) a ray-chaotic quarter bowtie billiard ($\mathcal{D}=2$), and (c) a three dimensional cavity with various symmetry breaking elements ($\mathcal{D}=3$). Scattering channels connected to the network analyzer (top right) are marked in red, and embedded perturbers are marked in green.
@@
The tunable perturbers make two types of measurements possible. One is a continuous parameter sweep like the kind shown in Fig. 1 where typically a single perturber is tuned with very fine steps between frequency scans, while the rest of the system parameters are held fixed. This kind of measurement reveals the phase winding around singularities, and through the use of a second perturber the dynamics of the singularities can be visualized [Erb2024, erb2025]. The other kind of measurement is a statistical ensemble, where each perturber is tuned simultaneously to very different values between frequency scans. This ensures that the difference between each realization in the ensemble is statistically significant. In other words, the first kind of measurement gives the scattering matrix in a closed area ($d=2$) or volume ($d=3$) of parameter space, while the second kind of measurement provides a reasonably ergodic sampling of a higher dimensional parameter space ($d\geq 4$). All the statistical results discussed in this paper will be derived from the second kind of measurement.
We consider four global system parameters as being important and hold these fixed across realizations of a given ensemble. These are: the dimension for wave propagation $\mathcal{D}$ ($\mathcal{D}\in\{1,2,3\}$), number of scattering channels $M$ ($M\in\{1,\dots,\infty\}$, although our network analyzer has a limit of $M\leq 4$), Dyson class $\beta$ ($\beta\in\{1,2\}$), and uniform cavity dissipation $\eta$ ($\eta\in\{0,\dots,\infty\}$). The Lorentz reciprocity of a system can be broken, changing $\beta$ from 1 to 2, by adding a magnetized ferrite in the propagation path of the microwaves [So1995, Stoffregen1995, Dietz2007, Dietz2009, Lawniczak2009, Lawniczak2010, Lawniczak2011, Bittner2014, Castaneda2022]. We approximate the cavity loss to be effectively homogeneous in all our experimental systems. In the $\mathcal{D}=1$ graph the absorption is due to dielectric loss and conductor resistivity in the coaxial cables and phase shifters, which are nominally uniform [Lei2022]. In the $\mathcal{D}=2$ billiard, the majority of the loss is caused by induced currents on the top and bottom plates which have uniform resistivity. Finally, the $\mathcal{D}=3$ cavity is highly reverberant, meaning the waves interact with the lossy boundaries many times.
@@
Note that $\eta$ is a dimensionless quantity given by $\eta:=\tau_{H}\tilde{\eta}$ where $\tilde{\eta}$ is an internal dissipation rate proportional to the average mode bandwidth [Buttiker1986, Brouwer1996, Kuhl2013]. The Heisenberg time $\tau_{H}=\frac{2\pi}{\Delta}$ can be related to the mean delay time of waves in a scattering environment, where $\Delta$ is the mean mode spacing in frequency units. Hence, the cavity absorption strength $\eta$ is inherently tied to our system dimension $\mathcal{D}$, as the larger systems have longer delay times which cause the waves to be more strongly attenuated. Since both $\tilde{\eta}$ and $\tau_{H}$ (for $\mathcal{D}>1$) are frequency dependent [Lei2022], by measuring the same system in different frequency bands we can systematically vary the value of $\eta$. Distributing absorbers in a cavity [Hemmady2005, Hemmady2005b, Hemmady2006, Hemmady2006c] or attenuators on the bonds of a graph [Hul2005, Lawniczak2023, Lawniczak2019] is another way to change $\eta$, though these methods can only increase the loss. One can also employ cryogenic cooling to decrease $\eta$ [Graf92, Hemmady2006, Dietz2015, Xiao2018, Ma2023]. For this paper we have prepared systems with $\eta$ varying from $1.8$ to $145$, which covers well what has been considered the low, moderate, and high absorption strength regimes [Sanchez2003, Hul2004, Hemmady2005, Fyodorov2004, Savin2005, Lawniczak2008, Lawniczak2019]. Table 1 in Appendix D has the values of the parameters and number of realizations in each measured ensemble.
@@
The $S$-matrix measured by a network analyzer is dependent on how well the scattering channels are coupled to the enclosure. Our systems are designed to be well coupled, but the coupling is still imperfect and frequency dependent. We remove these non-universal effects from ensembles through application of the Random Coupling Model (RCM) normalization process, which results in $S$-matrices with perfect coupling [Hemmady2005, Zheng20052Port, Zheng2006, Hart2009, Yeh2010, Lawniczak2012, Hemmady2012, Lee2013, Yeh2013, Gradoni2014, Auregan16, Dietz2017, Fu2017]. Coupling effects cannot be removed from parametric studies like in Fig. 1 because RCM requires statistical information. The effects of coupling on scattering singularity statistics are not addressed in this paper and will be explored in subsequent work. For more details on the experimental systems, such as more information on the tunable pertubers or how an ensemble value of $\eta$ is calculated, refer to Refs. [shaibe2025] and [erb2025].
@@ III Data and Discussion
In this section we present statistical results of scattering singularities from both experiment and RMT numerics. Note that every scattering matrix included is perfectly coupled, see Section II. We consider PDFs of functions that diverge at singularities for two reasons: (1) divergent functions are more advantageous for finding singularities, as compared to complex linear functions, because they require only one degree of freedom and have a more telling character when singularities occur, and (2) the PDF of a divergent function is dominated by an infinitely long tail which corresponds to the presence of singularities. This means most of the PDF is of importance when considering singularities, rather than just a small region around the singularity value. In particular, for every fundamental singularity, it is possible to define a time delay or some generalized delay that diverges at the singularity, as discussed in Appendix G. In subsection III.1 we present the statistics of topologically protected points (fundamental singularities) and curves in two dimensional parameter space. Subsection III.2 focuses on how absorption affects singularity abundance in lossy systems because absorption appears to be the most important determining parameter. Lastly, in subsection III.3 we discuss higher order singularities which show system-specific statistics, in contrast to fundamental singularities.
@@ III.1 Fundamental Singularities - Statistics
Figure 3: (a-b) PDFs of $\frac{\langle|\text{Re}[\delta R_{xy}]|\rangle}{|\text{Re}[\delta R_{xy}]|}$ and $\frac{\langle|\delta R_{xy}|\rangle}{|\delta R_{xy}|}$ from 31 ensembles of experimental scattering systems with different values for the four parameters $\mathcal{D},M,\beta,\eta$ (see Appendix D for an enumeration of the ensembles). Color gradient of curves corresponds to $\eta$ (not uniformly spaced) of the ensembles. Dashed black lines characterize the power law of the tail behavior. (c-d) PDFs of $\frac{\langle|\text{Re}[\mathcal{S}]|\rangle}{|\text{Re}[\mathcal{S}]|}$ and $\frac{\langle|\mathcal{S}|\rangle}{|\mathcal{S}|}$ for various functions $\mathcal{S}$ from two arbitrary ensembles. Dashed black lines characterize the power law of the tail behavior.
@@
In Fig. 3(a-b), we show the logarithmic PDFs of $\frac{\langle|\text{Re}[\delta R_{xy}]|\rangle}{|\text{Re}[\delta R_{xy}]|}$ and $\frac{\langle|\delta R_{xy}|\rangle}{|\delta R_{xy}|}$ from 31 ensembles of experimental scattering systems with different values for the four parameters $\mathcal{D},M,\beta,\eta$ (see Appendix D for the details and Table 1 which has the parameter values for each ensemble presented). Here , $\delta R_{xy}=S_{xx}-S_{yy}$ is the reflection difference between ports $x$ and $y$, and $\langle|\delta R_{xy}|\rangle$ is the ensemble mean of $|\delta R_{xy}|$. We normalize by the ensemble mean value to bring the PDFs together, as otherwise the PDFs from different systems can be distantly shifted horizontally from each other, making it difficult to compare the tails. We find that $\mathcal{P}(\frac{\langle|\text{Re}[\delta R_{xy}]|\rangle}{|\text{Re}[\delta R_{xy}|]})$ has a universal 2-2 power law tail that holds for all systems, represented by the dashed black line in panel (a), since $\frac{\langle|\text{Re}[\delta R_{xy}]|\rangle}{|\text{Re}[\delta R_{xy}]|}$ has a simple pole at $\text{Re}[\delta R_{xy}]=0$, a constraint on a single degree of freedom. Similarly, $\frac{\langle|\delta R_{xy}|\rangle}{|\delta R_{xy}|}$ has a simple pole at $\delta R_{xy}=0+i0$ which are points, specifically symmetric reflection points, and we find that $\mathcal{P}(\frac{\langle|\delta R_{xy}|\rangle}{|\delta R_{xy}|})$ has a universal 3-3 power law represented by the dashed line in panel (b).
@@
Three arbitrary choices were made in constructing Fig. 3(a-b). First, the choice of complex scalar function to investigate. There is nothing special about the reflection difference $\delta R_{xy}$, any function with singularities will produce the same statistical results. Second, in panel (a) we could have used the imaginary part of $\delta R_{xy}$ (as opposed to the real part) and the PDF tail behavior would have been the same; there is no particular reason to chose one over the other. Finally, plotting the PDFs of the magnitude, without regard to sign. This choice was made for simplicity as plotting both positive and negative values on a logarithmic scale is non-trivial. That was done, however, in Refs. [Lei2021Stats, shaibe2025] for the case of complex time delay, and there the same tails were seen on both the positive and negative sides of the distribution.
@@
In Fig. 3(c), we show the logarithmic PDFs of eight different functions which have simple poles at topologically protected curves: (i) $\frac{\langle|C|\rangle}{|C|}$, (ii) $\frac{\langle|\text{Re[det}S]|\rangle}{|\text{Re[det}S]|}$, (iii) $\frac{\langle|\text{Re}[S_{xx}]|\rangle}{|\text{Re}[S_{xx}]|}$, (iv) $\frac{\langle|\text{Re}[S_{xy}]|\rangle}{|\text{Re}[S_{xy}]|}$, (v) $\frac{\langle|\text{Re}[\delta R_{xy}]|\rangle}{|\text{Re}[\delta R_{xy}]|}$, (vi) $\frac{\langle|\text{Re}[\tau_{xx}]|\rangle}{|\text{Re}[\tau_{xx}]|}$, (vii) $\frac{\langle|\text{Re}[\tau_{xy}]|\rangle}{|\text{Re}[\tau_{xy}]|}$, and (viii) $\frac{\langle|\text{Re}[\tau_{WS}]|\rangle}{|\text{Re}[\tau_{WS}]|}$. Just as in panel (a), the same thing could be done using the PDFs of the imaginary parts of these functions rather than the real, except for $\frac{\langle|C|\rangle}{|C|}$ as $|C|$ is already a real number. All the data in panel (c) comes from one ensemble, a reciprocal two channel billiard ($\beta=1,M=2,\mathcal{D}=2$) with uniform absorption of $\eta=40$. Because the functions have simple poles at curves, the PDFs all have 2-2 power laws as represented by the dashed black curve. The deepest blue curve corresponding to $\mathcal{P}(\frac{\langle|C|\rangle}{|C|})$ has a downturn and transitions to a 3-3 power law. This can be explained by the fact that our experimental systems always have some small degree of non-reciprocity which is impossible to eliminate. Further, the network analyzer determines $S_{xy}$ and $S_{yx}$ independently and does not have arbitrary precision. For most purposes, this is a negligible effect but in non-reciprocal systems two-eigenvector orthogonality takes the dimension of points because it requires two conditions to be simultaneously satisfied (see Appendix C of Ref. [erb2025]), which induces the 3-3 power law. In RMT simulations that are perfectly reciprocal, this transition in the power law of the PDF tail from 2-2 to 3-3 is not seen (see Appendix E).
@@
Panel (d) of Fig. 3 has the logarithmic PDFs of ten different functions which have simple poles at topologically protected points: (i) $\frac{\langle|C|\rangle}{|C|}$ diverges at Orthogonality Points, (ii) $\frac{1-\langle|C|\rangle}{1-|C|}$ diverges at Exceptional Points, (iii) $\frac{\langle|\text{det}S|\rangle}{|\text{det}S|}$ diverges at Coherent Perfect Absortion, (iv) $\frac{\langle|S_{xx}|\rangle}{|S_{xx}|}$ diverges at Reflectionless Scattering Modes, (v) $\frac{\langle|S_{xy}|\rangle}{|S_{xy}|}$ diverges at Transmissionless Scattering Modes, (vi) $\frac{\langle|\delta R_{xy}|\rangle}{|\delta R_{xy}|}$ diverges at Symmetric Reflection Points, (vii) $\frac{\langle|\delta T_{xy}|\rangle}{|\delta T_{xy}|}$ diverges at Reciprocal Transmission Points, (viii) $\frac{\langle|\tau_{xx}|\rangle}{|\tau_{xx}|}$ diverges at Complex Reflection Time Delay zeros, (ix) $\frac{\langle|\tau_{xy}|\rangle}{|\tau_{xy}|}$ diverges at Complex Transmission Time Delay zeros, and (x) $\frac{\langle|\tau_{WS}|\rangle}{|\tau_{WS}|}$ diverges at Complex Wigner-Smith Time Delay zeros. All the data in this panel comes from an ensemble measurement of a non-reciprocal two channel graph ($\beta=2,M=2,\mathcal{D}=1$) with uniform absorption of $\eta=8$. Eigenvector orthogonality has dimension of d2d-2 in non-reciprocal systems, but dimension of d1d-1 in reciprocal systems, which is why $\mathcal{P}(\frac{\langle|C|\rangle}{|C|})$ appears in both panels (c) and (d). Unlike panel (c), there is no deviation from the expected 3-3 power law behavior for the tails in panel (d).
@@
Between panels (a) and (c), it should become clear that the PDF of every function in every system that diverges with a simple pole at a topologically protected curve has a 2-2 power law. Similarly, panels (b) and (d) convey that the PDF of every function in every system that diverges with a simple pole at a topologically protected point has a 3-3 power law. A very simple mathematical argument for why the universal power laws take these values is given in Appendix A. A figure showing the power laws of PDFs calculated from $S$-matrices constructed with RMT numerics is available in Appendix E. The fact that RMT numerics returns statistics for scattering singularities which is identical to our empirical results from microwave systems reveals that these statistics are not specific to any particular kind of wave scattering. Rather, these statistical results describe all generic scattering systems.
Notably, in Fig. 3 we do not present the PDFs of the inverse eigenvalue difference $\frac{1}{{|\lambda_{j}-\lambda_{k}|}}$ nor the inverse eigenvalue condition number $r^{-1}$. This is because neither of these functions have a simple pole at a singularity. Because of the eigenvalue repulsion mentioned before, the eigenvalue difference $|\lambda_{j}-\lambda_{k}|$ goes to zero non-linearly, hence $\frac{1}{{|\lambda_{j}-\lambda_{k}|}}$ is not a simple pole. Similarly, $r^{-1}$ actually has a branch point of the form $\frac{1}{\sqrt{(1-|C|)}}$ at an EP-2. Since they do not satisfy the condition of having a simple pole divergence only at a singularity, the PDFs of $\frac{1}{{|\lambda_{j}-\lambda_{k}|}}$ and $r^{-1}$ do not have 3-3 power laws. We provide some examples of PDFs of functions that diverge at singularities but not as simple poles in Appendix B.
@@
Since a distribution with a 3-3 power law has an undefined variance, a generic non-Hermitian scattering system must have every kind of scattering singularity allowed somewhere in its parameter space. However, it may not necessarily be accessible depending on the tunability of the experimental perturbers. Note also that since the 3-3 power law tail is “superuniversal”, the PDF of a function that diverges at a singularity is the perfect platform for investigating which global system parameters have the greatest effect on singularity abundance and density. It is also worth considering what the limitations of the “superuniversality” are, and whether there are ways to break it, for example by having imperfect coupling which was seen to disrupt the 3-3 power laws from some but not all complex time delay PDFs in Ref. [shaibe2025].